OpenStudy (anonymous):

f(x)=x+4 -3≤x<1 x=1 x>1 8 -x+2 DOMAIN? INTERCEPTS?

OpenStudy (anonymous):

domain is \[[-3,\infty)\] since the function is defined only for numbers greater than or equal to -3

OpenStudy (anonymous):

to find the y intercept set x = 0 and solve

OpenStudy (anonymous):

because \[-3\leq 0<1\] use the first rule and get \[f(0)=0+4=4\] so y intercept is (0,4)

OpenStudy (anonymous):

to find x intercepts set y = 0 and solve now \[x+4=0\implies x=-4\] but -4 is not in the domain of the function so forget about that one \[8=0\] is not possible \[-x+2=0\implies x=2\] so the x intercept is (2,0)